OpenStudy (anonymous):

Solve this system of equations using the addition method. x+y=12 x-y=4 @Mathematic Solve this system of equations using the addition method. x+y=12 x-y=4 @Mathematic @Mathematics

OpenStudy (hoblos):

x=8 y=4

hero (hero):

Add the two equations togther; y is eliminated. Solve for x: 2x = 16 x = 8 y=4

OpenStudy (anonymous):

o ok so like for x+y=3 and 2x-y+0 i add x+2x which gives me 3x then i add yand-y which gives me 0 and then 3 and 0 which is 3 so it would be 3x+0=3

OpenStudy (anonymous):

that looks wrong i did sumthin wrong

OpenStudy (anonymous):

what's x+x? what's y+(-y)? what's 12+4? 2x+0=16 2x=16 divide by 2. x=8 plug the x back into one of the starting equations.